# The Narrow Depth and Breadth of Corporate Responsible AI Research

### v1

@@ 4.1.5 Robustness of Classification Method
We tested in two different ways to validate whether the disparity observed above is robust to changes in the selection of dataset or methodology. First, we performed data ablation by varying the training and testing data. Second, we classified the papers with a different method.
@@
We also disaggregated the publications into conferences and journals (see Fig. S3a and Fig. S3b in the appendix, respectively). This analysis suggests that industry has a lower proportion of responsible AI research in both journal and conference participation; however, the disparity is more pronounced in journal participation.
@@ 4.1.6 Validating Results Using Expert-Suggested Keywords
To validate the findings from our ensemble model, we classified the papers using an alternate approach: by searching for expert-recommended keywords in the paper titles and abstracts. Here, we used a predefined list (i.e., dictionary) consisting of specific terms that indicated the category of interest, i.e., whether an observation matches the themes of responsible AI or not. However, a keyword-based search approach can be challenging in the absence of a relevant and representative dictionary. To address this concern, we utilized an extended list of keywords for responsible AI as suggested by experts in this field (see Table LABEL:table:oldA6 in the appendix). Similarly, when classifying conventional AI papers, we utilized keywords from the extant literature (Cockburn et al.,, 2019; Baruffaldi et al.,, 2020) and added additional words (see section S1.4 in the appendix).
@@
We used the same dataset employed in our prior analysis of industry engagement using the machine learning model: 679,919 papers between 2010 and 2022 co-authored by researchers affiliated with 1,771 leading AI firms, and 5,265,419 papers (2010-22) co-authored by researchers affiliated with 100 leading universities. Our findings from this keyword-based search on paper titles and abstracts, as shown in Fig. S7 in the appendix, align with our prior findings using the machine learning model and indicate that the majority of AI firms are engaging in responsible AI research to a very limited extent compared to academia. This methodological triangulation (Denzin,, 2017), employing varied approaches, evidently illustrates the disparity in responsible AI research between industry and academia.
@@ 4.1.7 Quality of Industry’s Responsible AI research
A common critique of industry involvement is the concern that its efforts may amount to mere superficial commitment rather than substantive (Vynck and Oremus,, 2023; Metcalf et al.,, 2019), rigorous research. In contrast, our analysis of the 50 most-cited papers from leading responsible AI conferences indicates that industry’s contribution to responsible AI research is not merely performative. Indeed, industry has made significant contributions in the form of high-quality, impactful research. We found that half of the 50 most-cited papers (2018-2022, leading three responsible AI papers) have authors with industry affiliations, and six of those papers are exclusively industry-authored. The large technology firms, Google and Microsoft, are the most represented, with ten and nine papers, respectively. Notably, nine of the 25 most-cited papers with industry affiliations involve authors who have since been dismissed or have left their organizations. Taken together, these results indicate that while industry has a narrower scope of responsible AI research than academia, it has nonetheless made a substantial contribution to the field in the past.
@@ 4.2 Assessing Industry Engagement in Responsible and Conventional AI Research
To assess the disparity in industry participation between responsible AI and conventional AI research, we conducted three distinct analyses using multiple datasets. First, we compared industry engagement in leading responsible AI conferences with its presence at the leading conventional AI conferences. This comparison was performed using two approaches: counting the unique number of firms and examining their co-authorship behavior. Second, we analyzed the growth of industry engagement in responsible AI research over time using a more comprehensive sample that extends beyond top-tier articles.
@@ 4.2.1 Data and Method
First, we selected three leading responsible AI conferences–ACM FAccT, AAAI/ACM AIES, and ACM EAAMO–based on our consultation with scholars researching responsible AI across multiple disciplines (computer science, information science, business, and philosophy). We found 851 papers in total from 2018 to 2022, of which 748 were downloaded from Scopus and 103 were downloaded from their respective websites (dataset 2). Then we manually classified all of these papers’ affiliations. We counted a paper as an industry paper if at least one of its co-authors had any industry affiliation, following prior literature (Frank et al.,, 2019). We then manually reviewed these papers’ affiliations to compile a list of unique firms represented in the authorship. Henceforth, this process of measuring industry engagement was applied separately for conferences and journals, allowing us to quantify and compare the level of industry engagement across different categories.
@@
Keywords-based distinct sample. This distinct sample is a key strength of our study, setting it apart from previous works (Maslej et al.,, 2024; Birhane et al.,, 2022) that focused solely on leading conferences. We adopted a more comprehensive definition of responsible AI, incorporating papers from a wider range of outlets beyond the top-tier conferences. This allows us to examine industry participation in less prominent outlets and provide a more comprehensive assessment.
@@ 4.2.2 Findings
First, we looked at the unique number of firms that have presented research at leading conventional AI conferences and responsible AI conferences. To measure this, first, we took the subset of papers where at least one of the co-authors was affiliated with industry. Then we manually curated the list of unique firms from the co-authors’ affiliations. Using that list, we show in Fig. 2a that only a small number of firms participate in responsible AI conferences. In case of conventional AI conferences (see Table S1 in the appendix for the list of conferences) more than 250 firms participated in each of all 10 of them; whereas, in stark contrast, the numbers for the responsible AI conferences (FAccT, AIES, and EAMMO) were significantly lower, with only 33, 28, and six firms participating, respectively (see Fig. 2a).
@@
Figure 2: Here, Fig. 2a shows a strong disparity in industry engagement by measuring firms’ presence (based on co-authors’ industry affiliation) in ten leading conventional AI conferences (n = 63,526 papers) and three leading responsible AI conferences (n = 851 papers) between 2018 and 2022. Fig. 2b shows the percentage of papers with at least one co-author having industry affiliation in conventional AI (n = 1,06,012; 2010-22), and responsible AI conferences (n = 851; 2018-22) over the years. This result is robust to changes in methods of co-authorship count. Fig. 2c represents the percentage of industry-papers (n = 36,022; 2010-22) classified as responsible AI papers (a distinct sample based on expert-suggested keywords) from both journals and conferences.
Subsequently, we examined industry co-authorship in the three major responsible AI conferences. Fig. 2b plots the proportion of industry co-authored papers in leading responsible AI conferences compared to the proportion of industry papers in conventional AI conferences (the dashed line). We calculated this annual proportion by dividing the total number of industry papers in each year for each conference by the total annual publications of each conference. The trends shown in Fig. 2b suggest that industry has a limited presence in responsible AI conferences while its participation is growing at conventional AI conferences. From 2018 to 2022, industry’s participation in responsible AI conferences has barely changed (the percentage varies roughly between 20% and 23%). [^6] Our supplementary analysis of ACM FAccT 2023 (not included in this study) reveals a downward trend in industry participation In contrast, during the same period, industry increased its participation from 29% to 37% at the leading conventional AI conferences. Our results are consistent with recent research (Ahmed et al.,, 2023) indicating industry has significantly increased its presence in conventional AI research.
Finally, using a separate sample based on expert-suggested keywords (dataset 3), we found that industry proportion has largely stayed the same in recent years (see Fig. 2c), which is consistent with our earlier findings. This suggests that, among all the peer-reviewed responsible AI publications over the years, annually, roughly 15% or less of them had a co-author affiliated with industry. Even more worryingly, this number has largely stayed the same in recent years. To conclude, even when we used a different sample, we found that industry presence in responsible AI research has been limited.
Overall, our results illustrate the divergence between industry participation in conventional AI and responsible AI research. Different interpretations of the criteria defining an industry paper produce similar outcomes.
@@ 4.2.3 Alternative Measurements
Responsible AI research at leading conventional AI conferences: Recent research indicates the growing presence of responsible AI research at conventional AI conferences (Maslej et al.,, 2024). This trend raises the question of whether industry is more focused on engaging in responsible AI research at these conferences. To investigate that possibility, we analyzed industry’s contribution to responsible AI research presented at leading conventional AI conferences. Our findings reveal that industry’s share of responsible AI research at these conferences has largely stayed the same in recent years (see Fig. S8 in the appendix).
@@ 4.3 Summary of Findings
In this section, by analyzing nearly 6 million papers from industry and academia, we showed that industry lacks meaningful engagement in responsible AI research, in both quantity and impact, compared to academia. Additionally, among 519 firms conducting AI research, only about 11.2% of them meaningfully engaged in responsible AI research through peer-reviewed publications. By examining industry presence in leading conferences on conventional and responsible AI, we also observe industry’s noticeably lower presence in responsible AI conferences. While industry’s participation in responsible AI conferences has remained fairly the same, we showed that their presence in conventional AI conferences has increased over the years in contrast. Overall, our analysis, including calculating the number of co-authoring firms and the percentage of industry papers, shows industry’s limited presence in responsible AI research.
@@ 5.2 K-means Clustering
To analyze the research priorities of industry and academia separately in responsible AI research, we used k-means clustering (MacQueen,, 1967) in our data. K-means clustering is used to organize large volumes of unstructured text data into meaningful clusters, providing a foundation for an intuitive understanding.
@@
This clustering, shown in Fig. 3a, highlights the narrow breadth of industry’s responsible AI research. Here we quantify how differently industry and academia prioritize different issues in responsible AI research. Our results suggest that industry focuses mostly on issues like “bias,” “explainable AI,” “human-AI interaction,” and “algorithmic fairness,” often sidelining broader societal implications. Academia, along with “bias,” emphasizes more on “human-centric AI,” “ethical and moral concerns,” and “equitable AI.” Both sectors, however, converge on the importance of “privacy” and “data governance” (see Table LABEL:table:3AClusters in the appendix for associated terms).
We conducted sensitivity analyses by varying the number of clusters, and our findings remained consistent (see Fig. S12 and Fig. S13 in the appendix). Additionally, our temporal analysis suggests industry’s focus on responsible AI research gets narrower over time, and industry tends to engage in certain topics only after academia introduces them (see Fig. S9 in the appendix).
A comparison with conventional AI conferences. We compared the research priorities of industry and academia by examining their focus in both conventional AI and responsible AI conferences. We performed separate clustering analyses on industry-only and academia-only papers from the leading ten conventional AI conferences (see Fig. S11 in the appendix). Interestingly, we found that the research focus of industry and academia in these conventional AI conferences are largely similar. This observation highlights industry’s limited engagement in responsible AI research, as their research focus in conventional AI conferences closely mirrors that of academia, while their research breadth in responsible AI remains narrower.
@@ 5.3 Structural Topic Modeling
Next, we used structural topic modeling (STM) (Roberts et al.,, 2013) to identify how prevalent different research topics are across industry and academia. STM builds on probabilistic topic modeling (Blei,, 2012) and is widely used in the literature (Lucas et al.,, 2015; Mulder et al.,, 2021; Laufer et al.,, 2022). This method allows additional information, such as affiliations, to be incorporated along with the target text to be analyzed to estimate topic proportions more precisely. This let us analyze how prevalent (or dominant) a specific topic was across our target groups (i.e., industry and academia). In our analysis, we used the papers’ authors’ affiliations (i.e., industry or academia) as additional metadata along with abstracts. We used the same dataset here as we did in our earlier k-means clustering analysis.
@@
Figure 3: Here, the first figure Fig. 3a quantifies the key focus areas in responsible AI research between industry and academia (n = 10,799 papers in total; 2010-22) using k-means clustering on paper abstracts. Fig. 3b shows the structural topic modeling estimates using the same data. Topics toward the right-hand side are more prevalent in abstracts from industry, while those toward the left are more prevalent in abstracts from academia. Both analyses show that industry, unlike academia, prioritizes more in technical fixes, than in ethical and moral concerns, and societal implications. Subsequently, by conducting a frequency analysis of relevant terms on the same data, Fig. 3c shows the percentage that industry has limited engagement in issues like, “Human rights” and “Environmental concerns,” while emphasizing more on “Explainable AI.”
@@
Our STM analysis, by analyzing the dominant topics present in both groups’ work, suggests a concerning difference between the research priorities of industry and academia. This shows that industry heavily emphasizes research on developing and improving AI technology. In contrast, they pay limited attention to moral and ethical issues, accountability, and the societal implications of their deployed technologies. Even though there is a growing recognition within computing research that computing tools have an important role to play in addressing social issues (Abebe et al.,, 2020), firms’ engagement in responsible AI research seems to overlook such normative issues. Overall, our results suggest that industry is more focused on technical rather than sociotechnical aspects of responsible AI.

### v2

@@ 4.1.5 Robustness of Classification Method
We tested in three different ways to validate whether the patterns observed above are robust to changes in the selection of dataset or methodology. First, we performed data ablation by varying the training and testing data. Second, we classified the papers with a different method.
@@
We also disaggregated the publications into conferences and journals (see Fig. S3a and Fig. S3b in the appendix, respectively). This analysis suggests that industry has a lower proportion of responsible AI research in both journal and conference publications; however, the difference is larger in journal publications.
@@ 4.1.6 Validating Results Using Expert-Suggested Keywords
To validate the findings from our ensemble model, we classified the papers using an alternate approach: by searching for expert-recommended keywords in the paper titles and abstracts. Here, we used a predefined list (i.e., dictionary) consisting of specific terms that indicated the category of interest, i.e., whether an observation matches the themes of responsible AI or not. However, a keyword-based search approach can be challenging in the absence of a relevant and representative dictionary. To address this concern, we utilized an extended list of keywords for responsible AI as suggested by experts in this field (see Table LABEL:table:oldA6 in the appendix). Similarly, when classifying conventional AI papers, we utilized keywords from the extant literature (Cockburn, Henderson, and Stern, 2019; Baruffaldi, van Beuzekom, Dernis, Harhoff, Rao, Rosenfeld, and Squicciarini, 2020) and added additional words (see section S1.5 in the appendix).
@@
We used the same dataset employed in our prior analysis of industry output using the machine learning model: 679,919 papers between 2010 and 2022 co-authored by researchers affiliated with 1,771 leading AI firms, and 5,265,419 papers (2010-22) co-authored by researchers affiliated with 100 leading universities. Our findings from this keyword-based search on paper titles and abstracts, as shown in Fig. S7 in the appendix, align with our prior findings using the machine learning model and indicate that AI firms exhibit lower rates of responsible AI research output relative to academia. This methodological triangulation (Denzin, 2017), employing varied approaches, yields consistent results regarding the difference in responsible AI research between industry and academia.
4.1.7 Validating Results Using LLM-based Classification
To assess the robustness of our findings, we employed an alternative classification approach using large language models (LLMs) with in-context learning (ICL). This method leverages an LLM’s capacity to perform classification tasks based on annotated examples without fine-tuning model parameters (Xu, Cohen, Wang, and Srikumar, 2024; Tay, Tran, Dehghani, Ni, Bahri, Mehta, Qin, Hui, Zhao, Gupta, Schuster, Cohen, and Metzler, 2022). We evaluated multiple proprietary LLMs, including Gemini 1.5 Flash, GPT-4o, OpenAI o1, and Claude 3.5 Sonnet (versions as of October 2024). For each model, we prepared sets of 47 and 44 manually annotated samples to classify papers as conventional AI, responsible AI, or non-AI. We assessed model performance through data ablation across multiple hold-out subsets and validated results both manually and against our ensemble model. On the basis of this evaluation, we selected Gemini 1.5 Flash for subsequent validation. Finally, using LLM as a classifier produced results consistent with our prior analyses. As shown in Fig. S8 in the appendix, the patterns of industry and academic engagement in responsible AI research observed using LLM classification correspond with those obtained from both our supervised machine learning model and keyword-based approach. This convergence across three distinct classification methods confirms the reliability of our findings.
@@ 4.1.8 Quality of Industry’s Responsible AI research
Prior literature has raised questions about the nature and depth of industry involvement in responsible AI research (Vynck and Oremus, 2023; Metcalf, Moss, and Boyd, 2019). To examine this dimension, we analyzed the 50 most-cited papers from leading responsible AI conferences. We found that half of these highly cited papers (2018–2022, leading three responsible AI conferences) have authors with industry affiliations, and six are exclusively industry-authored. This substantial representation among the field’s most influential work suggests that industry researchers have contributed meaningfully to shaping the responsible AI research agenda. Google and Microsoft are the most represented firms, with ten and nine papers, respectively, reflecting sustained institutional investment in this area. Several other firms, including IBM, Meta, and DeepMind, also appear among these highly cited contributions. These findings indicate that while industry’s responsible AI output is lower in volume than academia’s, the work produced by active industry researchers has achieved considerable scholarly impact and recognition within the field.
@@ 4.2 Assessing Industry Engagement in Responsible and Conventional AI Research
To assess the variation in industry participation between responsible AI and conventional AI research, we conducted three distinct analyses using multiple datasets. First, we compared industry engagement in leading responsible AI conferences with its presence at the leading conventional AI conferences. This comparison was performed using two approaches: counting the unique number of firms and examining their co-authorship behavior. Second, we analyzed the growth of industry engagement in responsible AI research over time using a more comprehensive sample that extends beyond top-tier articles.
@@ 4.2.1 Data and Method
First, we selected three leading responsible AI conferences–ACM FAccT, AAAI/ACM AIES, and ACM EAAMO–based on our consultation with scholars researching responsible AI across multiple disciplines (computer science, information science, business, and philosophy). We found 851 papers in total from 2018 to 2022, of which 748 were downloaded from Scopus and 103 were downloaded from their respective websites (dataset 2). Then we manually classified all of these papers’ affiliations. We counted a paper as an industry paper if at least one of its co-authors had any industry affiliation, following prior literature (Frank, Wang, Cebrian, and Rahwan, 2019). We then manually reviewed these papers’ affiliations to compile a list of unique firms represented in the authorship. Henceforth, this process of measuring industry engagement was applied separately for conferences and journals, allowing us to quantify and compare the level of industry engagement across different categories.
@@
Keywords-based distinct sample. This distinct sample is a key strength of our study, setting it apart from previous works (Maslej, Fattorini, Perrault, Parli, Reuel, Brynjolfsson, Etchemendy, Ligett, Lyons, Manyika, Niebles, Shoham, Wald, and Clark, 2024; Birhane, Kalluri, Card, Agnew, Dotan, and Bao, 2022) that focused solely on leading conferences. We adopted a more comprehensive definition of responsible AI, incorporating papers from a wider range of outlets beyond the top-tier conferences. This allows us to examine industry participation in less prominent outlets and provide a more comprehensive assessment.
@@ 4.2.2 Findings
First, we examined the unique number of firms that have presented research at leading conventional AI conferences and responsible AI conferences. To measure this, we took the subset of papers where at least one co-author was affiliated with industry and manually curated the list of unique firms from co-authors’ affiliations. Fig. 2a presents the distribution of firm participation across these venues. For conventional AI conferences (see Table S1 in the appendix for the list of conferences), more than 250 firms participated in each of all 10 venues. In contrast, participation at responsible AI conferences (FAccT, AIES, and EAAMO) was more concentrated, with 33, 28, and six firms participating, respectively (see Fig. 2a). This variation may reflect differences in conference size, disciplinary focus, or the relative maturity of these venues within the broader AI research ecosystem.
@@
Figure 2: Here, Fig. 2a shows the distribution of industry participation by measuring firms’ presence (based on co-authors’ industry affiliation) in ten leading conventional AI conferences (n = 63,526 papers) and three leading responsible AI conferences (n = 851 papers) between 2018 and 2022. Fig. 2b shows the percentage of papers with at least one co-author having industry affiliation in conventional AI (n = 106,012; 2010-22), and responsible AI conferences (n = 851; 2018-22) over the years. This result is robust to changes in methods of co-authorship count. Fig. 2c represents the percentage of industry-papers (n = 36,022; 2010-22) classified as responsible AI papers (a distinct sample based on expert-suggested keywords) from both journals and conferences.
Subsequently, we examined industry co-authorship in the three major responsible AI conferences. Fig. 2b plots the proportion of industry co-authored papers in leading responsible AI conferences compared to the proportion of industry papers in conventional AI conferences (the dashed line). We calculated this annual proportion by dividing the total number of industry papers in each year for each conference by the total annual publications of each conference. The trends shown in Fig. 2b suggests that industry participation rates differ between responsible AI conferences and conventional AI conferences, with participation growing in the latter. From 2018 to 2022, industry’s participation in responsible AI conferences remained relatively stable (the percentage varies roughly between 20% and 23%). [^6] Our supplementary analysis of ACM FAccT 2023 (not included in this study) reveals a similar trend in industry participation. By comparison, during the same period, industry increased its participation from 29% to 37% at the leading conventional AI conferences. Our results are consistent with recent research (Ahmed, Wahed, and Thompson, 2023) indicating industry has increased its presence in conventional AI research.
Finally, using a separate sample based on expert-suggested keywords (dataset 3), we found that industry proportion has largely stayed the same in recent years (see Fig. 2c), which is consistent with our earlier findings. This suggests that, among all the peer-reviewed responsible AI publications over the years, annually, roughly 15% or less of them had a co-author affiliated with industry. This proportion has remained stable over the observation period. These patterns are consistent across different samples and classification methods.
Overall, our results document different trajectories in industry participation between conventional AI and responsible AI research. Different interpretations of the criteria defining an industry paper produce similar outcomes.
@@ 4.2.3 Alternative Measurements
Responsible AI research at leading conventional AI conferences: Recent research indicates the growing presence of responsible AI research at conventional AI conferences (Maslej, Fattorini, Perrault, Parli, Reuel, Brynjolfsson, Etchemendy, Ligett, Lyons, Manyika, Niebles, Shoham, Wald, and Clark, 2024). This trend raises the question of whether industry is more focused on engaging in responsible AI research at these conferences. To investigate that possibility, we analyzed industry’s contribution to responsible AI research presented at leading conventional AI conferences. Our findings reveal that industry’s share of responsible AI research at these conferences has largely stayed the same in recent years (see Fig. S9 in the appendix).
@@ 4.3 Summary of Findings
In this section, by analyzing nearly 6 million papers from industry and academia, we documented patterns of industry engagement in responsible AI research. Comparing leading AI firms with leading academic institutions, we observe that industry’s responsible AI research output is lower in both quantity and impact relative to academia, though considerable heterogeneity exists across firms, with several organizations demonstrating publication rates comparable to elite research universities. By examining industry presence in leading conferences on conventional and responsible AI, we observe that industry participation rates are lower in responsible AI conferences than in conventional AI conferences. While industry’s participation in responsible AI conferences has remained relatively stable, their presence in conventional AI conferences has increased over the years. Additionally, analysis of highly cited papers indicates that industry researchers have contributed substantially to influential work in the field. Overall, our analysis shows that responsible AI research constitutes a smaller share of industry’s research portfolio relative to academia’s, though the quality and impact of active industry contributors is recognized within the scholarly community.
@@ 5.2 K-means Clustering
To analyze the research themes of industry and academia separately in responsible AI research, we used k-means clustering (MacQueen, 1967) on our data. K-means clustering is used to organize large volumes of unstructured text data into meaningful clusters, providing a foundation for an intuitive understanding.
@@
This clustering, shown in Fig. 3a, illustrates the topical distribution of industry’s responsible AI research. Here we quantify the distribution of industry and academia research across different topics in responsible AI research. Our results suggest that industry research concentrates on issues like bias,” explainable AI,” human-AI interaction,” and algorithmic fairness,” with less representation in topics addressing broader societal implications. Academia, along with bias,” shows greater concentration in human-centric AI,” ethical and moral concerns,” and equitable AI.” Both sectors, however, converge on privacy” and data governance” (see Table LABEL:table:3AClusters in the appendix for associated terms).
We conducted sensitivity analyses by varying the number of clusters, and our findings remained consistent (see Fig. S13 and Fig. S14 in the appendix). Additionally, our temporal analysis suggests industry’s thematic distribution has become more concentrated over time, and industry tends to publish on certain topics after academia introduces them (see Fig. S10 in the appendix).
Comparison with conventional AI conferences. We compared the research themes of industry and academia by examining their distribution in both conventional AI and responsible AI conferences. We performed separate clustering analyses on industry-only and academia-only papers from the leading ten conventional AI conferences (see Fig. S12 in the appendix). We found that the research themes of industry and academia in these conventional AI conferences are largely similar. In contrast, the thematic distribution differs between industry and academia in responsible AI research, with industry research distributed across fewer topic clusters than academia.
@@ 5.3 Structural Topic Modeling
Next, we used structural topic modeling (STM) (Roberts, Stewart, Tingley, and Airoldi, 2013) to identify how different research topics are distributed across industry and academia. STM builds on probabilistic topic modeling (Blei, 2012) and is widely used in the literature (Lucas, Nielsen, Roberts, Stewart, Storer, and Tingley, 2015; Mulder, Inel, Oosterman, and Tintarev, 2021; Laufer, Jain, Cooper, Kleinberg, and Heidari, 2022). This method allows additional information, such as affiliations, to be incorporated along with the target text to be analyzed to estimate topic proportions more precisely. This let us analyze the relative prevalence of specific topics across our target groups (i.e., industry and academia). In our analysis, we used the papers’ authors’ affiliations (i.e., industry or academia) as additional metadata along with abstracts. We used the same dataset here as we did in our earlier k-means clustering analysis.
@@
Figure 3: Here, the first figure Fig. 3a quantifies the key focus areas in responsible AI research between industry and academia (n = 10,799 papers in total; 2010-22) using k-means clustering on paper abstracts. Fig. 3b shows the structural topic modeling estimates using the same data. Topics toward the right-hand side are more prevalent in abstracts from industry, while those toward the left are more prevalent in abstracts from academia. Both analyses indicate that industry research is more concentrated in technical topics, while academia shows greater concentration in ethical, moral, and societal topics. Subsequently, by conducting a frequency analysis of relevant terms on the same data, Fig. 3c shows the relative frequency of topic-related terms, with industry research showing lower frequencies for terms related to “Human rights” and “Environmental concerns,” and higher frequencies for terms related to “Explainable AI”.
@@
Our STM analysis, by analyzing the dominant topics present in both groups’ work, suggests a concerning difference between the research priorities of industry and academia. This shows that industry heavily emphasizes research on developing and improving AI technology. In contrast, they pay limited attention to moral and ethical issues, accountability, and the societal implications of their deployed technologies. Even though there is a growing recognition within computing research that computing tools have an important role to play in addressing social issues (Abebe, Barocas, Kleinberg, Levy, Raghavan, and Robinson, 2020), firms’ engagement in responsible AI research seems to overlook such normative issues. Overall, our results suggest that industry is more focused on technical rather than sociotechnical aspects of responsible AI.